OpenStudy (thcsbabygirl1224):

help me please 4x+y+5z=-9 x-4y-2x+-2 2x+3y-2z=21

OpenStudy (anonymous):

Use matrix

OpenStudy (thcsbabygirl1224):

what its a system of equations solving for 3 variables so im lost

OpenStudy (thcsbabygirl1224):

@hartnn

OpenStudy (thcsbabygirl1224):

@LazyBoy @hba

OpenStudy (thcsbabygirl1224):

@chainedecho

OpenStudy (loser66):

The second equation should be x-4y-2z =-2, right?

OpenStudy (thcsbabygirl1224):

yes

OpenStudy (loser66):

take the third one subtract the second one, what do you get? that is (3) -(2) =??

OpenStudy (thcsbabygirl1224):

i am trying to get rid of the x in all 3 equations so i have to set them to 0 im lost on the work

OpenStudy (loser66):

YOu lost on your way, give it a try on my way, ok? Again (equation 3 )-(equation 2) is??? You should work on it, I am guiding you only.

OpenStudy (thcsbabygirl1224):

@kal1921 what is the matrics

OpenStudy (loser66):

Good luck.

OpenStudy (thcsbabygirl1224):

ok i have now -2x-4y-2z=-4 2x+3y-2z=21

OpenStudy (thcsbabygirl1224):

@Michele_Laino

OpenStudy (michele_laino):

here, you have a system of two equations, with three unknows. We can rewrite it as below: \[\left\{ \begin{gathered} 2x + 4y = 4 - 2z \hfill \\ 2x + 3y = 21 + 2z \hfill \\ \end{gathered} \right.\] which can, easily solved for x, and y as functions of z. Please try

OpenStudy (thcsbabygirl1224):

You get 7y-4z=16